An all-equity firm is considering the projects shown below. The T-bill rate is 6 percent and the market risk premium is 8 percent.

Project -  Expected - Return Beta

A   9 % 0.3

B 21 % 1.1

C 15 % 1.3

D 19 % 1.5

Calculate the project-specific benchmarks for each project. (Round your answers to 2 decimal places.)
Answer #1

Benchmark for Project A = T-bill rate + beta of A * market risk premium = 6% + 0.3 * 8% = 8.40%

Benchmark for Project B = T-bill rate + beta of B * market risk premium = 6% + 1.1 * 8% = 14.80%

Benchmark for Project C = T-bill rate + beta of C * market risk premium = 6% + 1.3 * 8% = 16.40%

Benchmark for Project D = T-bill rate + beta of A * market risk premium = 6% + 1.5 * 8% = 18.00%